the time taken to clean up a particular fast food outlet after it closes follows a normal distribution with a mean of 30 minutes and a standard deviation of 5 minutes.

a. what is the probability that the cleanup crew will complete the job in less than 20 minutes?
b. What is the probability that the cleanup crew will complete the job in more than 38 minutes?
1
Expert's answer
2019-11-11T13:31:43-0500

Let "X=" the time taken to clean up a particular fast food outlet after it closes in minutes:"X\sim N(\mu, \sigma^2)"

Then


"Z={X-\mu \over \sigma}\sim N(0,1)"

Given that "\mu=30\ minutes, \sigma=5 \ minutes"

a.


"P(X<20)=P(Z<{20-30 \over 5})=P(Z<-2)\approx"

"\approx0.022750"

 The probability that the cleanup crew will complete the job in less than 20 minutes is "0.0228"


b.


"P(X>38)=1-P(X\leq38)=1-P(Z\leq{38-30 \over 5})="

"=1-P(Z\leq1.6)\approx1-0.945201\approx0.0548"

The probability that the cleanup crew will complete the job in more than 38 minutes is "0.0548"